OpenStudy (anonymous):

Solve 5^x + 5 = 9^x Help please..? @Zarkon

OpenStudy (zarkon):

you have to solve it numerically

OpenStudy (anonymous):

Here is the first step... \[x+5 \log5=x \log 9\] Then you use the calculator to find the logs of 9 and 5. That's where I get stuck.

OpenStudy (zarkon):

you will have to solve that numerically too,

OpenStudy (anonymous):

Ok, how would I do that?

OpenStudy (zarkon):

use a calculator or computer program that has a solve function

OpenStudy (mrnood):

\[x(\log(9)-1)= 5\log5\]

OpenStudy (mrnood):

so re-arrange for x and solve with calculator

OpenStudy (zarkon):

oops...or do that...for some reason when I wrote is down to do I put 5log(x) ;)

OpenStudy (anonymous):

So you find what the log of 9 and 5 is then you distribute that to the coefficients?

OpenStudy (mrnood):

Hold on - I think golf4lifes's formula may be wrong...

hero (hero):

@Golf4Life, the way you attempted to isolate x is not correct.

OpenStudy (mrnood):

I agree - hold on there....

OpenStudy (anonymous):

ok, so log5= 0.6989700x+ 5.6989700=log9=0.9542425x ..How do you simplify from here on?

OpenStudy (mrnood):

if you take logs of both sides you have log(5^x+5) i.e. log(a+b) Now, this is NOT equal to loga + logb, and is not generally able to be simplified (afaik) So - can I ask you to check the question: Is it \[5^{(x+5)}\] or \[5^{x}+5\]

OpenStudy (anonymous):

Yes it is the first one, 5^(x+5) @MrNood

OpenStudy (mrnood):

Ok - that makes quite a difference \[5^{x+5}=9^{x}\] \[(x+5)\log(5)=xlog(9)\] \[xlog(5)+5\log(5)=xlog(9)\] \[x(\log(5)-\log(9))=-5\log(5)\]

OpenStudy (anonymous):

is it -3.63?

hero (hero):

If that's what it is then: \[5^{x + 5} = 9^x\] \[5^x5^5 = 9^x\] \[\frac{5^x}{9^x}\dot\ 5^5 = 1\] \[\frac{5^x}{9^x} = \frac{1}{5^5}\] \[\left(\frac{5}{9}\right)^x = \frac{1}{5^5}\] \[\log{\left(\frac{5}{9}\right)^x} = \log\left({\frac{1}{5^5}}\right)\] \[x\log{\left(\frac{5}{9}\right)} = \log\left({\frac{1}{5^5}}\right)\] \[x = \frac{\log(1/5^5)}{\log(5/9)}\] \[x = \frac{\log(1) - 5\log(5)}{\log{5} - \log{9}}\] \[x = -\frac{5\log(5)}{\log{5} - \log{9}}\]

OpenStudy (mrnood):

log9>log5 so LHS is -ve RHS is -ve so answer must be +ve

OpenStudy (mrnood):

be sure to notice the minus sign on Hero's equation - otherwise it is the same as my approach

OpenStudy (mrnood):

?

OpenStudy (anonymous):

x ≈ -3.63 x ≈ -0.28 x ≈ 0.07 x ≈ 13.69 These are the choices. When I put Hero's equation into the calculator I get -4.05

OpenStudy (anonymous):

@Hero

hero (hero):

When I put my answer into my calculator, I get one of the answer choices.

hero (hero):

@Golf4Life, I expressed x in three different ways. You should test them all out in your calculator.

hero (hero):

Each result should be the same.

OpenStudy (anonymous):

I put each of your problems in, maybe I am entering them wrong.\[x = \frac{\log(1) - 5\log(5)}{\log{5} - \log{9}}\] So for this one the top would be \[x=-3.49485\-0.25527\] Is that right? That should equal=-3.75 Which is choice (a)?

hero (hero):

Yep, you're confused about something.

hero (hero):

What kind of calculator do you have?

OpenStudy (anonymous):

I was using my iPhone. Should I use my Texas instrument one instead?

hero (hero):

Uh...ya. Your iPhone is not a replacement for your TI

hero (hero):

As a matter of fact, I use a TI Nspire. I recommend you get one as well.

OpenStudy (anonymous):

I have the TI-30XA. So how I put the equation in was it correct?

hero (hero):

If you didn't get the correct answer, then no, it isn't correct.

OpenStudy (anonymous):

Which is the correct one then?

OpenStudy (anonymous):

On both calculators I am getting the same thing:/

OpenStudy (anonymous):

@Hero?

hero (hero):

Probably because you're entering the same incorrect information into both calculators. I remember when I was your age. I used to do the same thing. It's a phase. You'll grow out of it (hopefully).

hero (hero):

Hang on a minute...let me help you out a bit.

hero (hero):

This is how you enter it into your calculator: -(5 log(5))/(log(5) - log(9))

hero (hero):

It has to be entered exactly like that or you won't get the right answer.

hero (hero):

Here's the other form I presented x: http://www.wolframalpha.com/input/?i=+log%281%2F5%5E5%29%2F%28log%285%2F9%29 And I entered: log(1/5^5)/(log(5/9) I got the same result again

OpenStudy (anonymous):

You are right! I put it in the way you said and I got 13.69. Thanks @Hero! How did you get so good at this stuff? :)

hero (hero):

The more familiar you become with something, the more comfortable you are with it.

OpenStudy (anonymous):

Haha ok thanks again!